Is is believed that in a certain town 35% of the people did not graduate from high school
a. If we sample 18 yr old or older, what is the probability the first or the second we talk to did not graduate high school?
b. If we samole adults 18 yrs or older what is the probability the third person sampled will be the first one who did not graduate high school?

Answer by gonzo(654)

not terribly sure, but reasonably sure that this is what they are looking for.
-----
first part
-----
probability that first or second you talk to didn't graduate from high school equals 1 minus probability that both that you talked to did graduate from high school.
-----
if p(did not graduate) = .35, then p(did graduate) = 1 - .35 = .65
since p(did graduate) = .65, then probability of first 2 persons graduating high school = .65*.65 = .4225.
if this is true, then the p(either didn't graduate) = 1 - p(both graduated) = 1 - .4225 = .5775
----
second part
-----
probability that the third person you interviewed is the first person that graduated from high school equals the probability that the first two didn't graduate from high school multiplied by the probability that the 3d person did graduate.
-----
probability that the first 2 persons you talked to didn't graduate from high school is p(didn't graduate) * p (didn't graduate) = .35*.35 = .1225.
p(first 2 persons didn't graduate) = .1225
probability that the third person did graduate is .65.
probability that the first 2 persons didn't graduate and the third person did graduate is
p(first 2 persons didn't graduate) * p(third person graduated) = .1225*.65 = .079625.
-----
answer for part 1 is .5775
answer for part 2 is .079625
-----
if i assumed correctly.
i believe i did.
-----
i'm not totally sure but i think this removes your question from the list of unsolved problems.
if it does, and you want to get further answers, then resubmit.
check to see if it's still there as unsolved.
if it is or isn't, please let me know.
thanks